OpenStudy (anonymous):

I AM HELPING MY SIS OUT AND SHE DONT UNDERSTAND PLAY HELP A gift box is 12 inches long, 8 inches wide, and 2 inches high. How much wrapping paper is needed to wrap the gift box? a)272 in2 b)192 in2 c)384 in2 d)184 in2 18. Tamara needs to buy motor oil to fill the 3 empty cylindrical barrels at her oil service center. Each barrel is 7 ft deep and has a radius of 4 ft. What is the volume of oil needed? Use = 3.14. a)1,055.04 ft3 b)4,220.16 ft3 c)602.88 ft3 d)1,406.72 ft3

OpenStudy (anonymous):

LOL PLAY I MEN PLZ

OpenStudy (anonymous):

Find the surface area of the object for number one.

OpenStudy (anonymous):

Find the volume of the barrel for number two.

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

AND

OpenStudy (anonymous):

HELLLLOOOO

OpenStudy (anonymous):

Hi, give me a sec to get the right formulas. brb

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

#2 is V = (pi)(r^2)(h) V = (3.14)(4^2)(7) V = (3.14)(16)(7) V = (351.68)(3) V = 1055.04 The answer is A

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

#1 is 2lw + 2lh + 2wh l = 12 w = 8 h = 2 2(12)(8)+2(12)(2)+2(8)(2) 2(96)+2(24)+2(16) 192+48+32 272 The answer is A also.

OpenStudy (anonymous):

WOW THX

OpenStudy (anonymous):

NP

OpenStudy (anonymous):

ONE MORE PLZZZ

OpenStudy (anonymous):

Find the volume of a triangular prism with the following dimensions. Round to the nearest whole number. triangle base: 5 ft triangle height: 3 ft prism height: 7.5 ft (1 point) a)56 ft3 b)19 ft3 c)113 ft3 D)8 ft3

OpenStudy (anonymous):

ok,give me a sec though

OpenStudy (anonymous):

OK TYT

OpenStudy (anonymous):

Working on it now…

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

V = Area of base x height A = 1/2 (b)(h1) A = ½(5)(3) A = ½(15) A = 7.5 V = (7.5)(7.5) V = 56.25 V = 56 (rounded) The answer is A... I see a pattern

OpenStudy (anonymous):

WOW YOUR GOOD

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

That it?

OpenStudy (anonymous):

NOT REALLY BUT YOU DONT HAVE TO HELP IF YOU DONT WHAT TO

OpenStudy (anonymous):

umm.... fan me an' I'll do 1 more

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

20. When baking a cake, you have a choice of the following pans: a round cake pan that is 2 inches deep and has a 7 inch diameter a 6 inch × 9 inch rectangular cake pan that is 2 inches deep a. Which of these pans has the larger volume? Justify your answer. b. How much icing will be needed to frost each cake, excluding the bottom of the cake? Justify your answer.

OpenStudy (anonymous):

essay huh? ok, give me a sec

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

A. Pan #1 is a cylinder, Pan #2 is a rectangular prism. Pan #1: V = (3.14)(r^2)(h) V = (3.14)(3.5^2)(2) V = (3.14)(12.25)(2) V = 76.93 in^3 V = 77 in^3 (rounded) Pan #2: V = (l)(w)(h) V = (6)(9)(2) V = (6)(18) V = 108 in^3 Pan #2 has a higher volume. Gimme a sec for part 2.

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

YOU THEIR

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

Part 2 is a surface area minus a face. Just calculate the area of the face and set it aside. Calculate the surface area of that shape and subtract the area of the face. #1: Area of a circle is: A = (pi)(r^2) A = (3.14)(3.5^2) A = (3.14)(12.25) A = 38.47 Surface area of a cylinder: SA = 2 pi r 2 + 2 pi r h 2(3.14)(3.5^2) + 2(3.14)(3.5)(2) 2(3.14)(12.25) + 2(3.14)(3.5)(2) 2(38.47) + 2(22) 77 + 44 121 121-38.47 = 82.53 82.53 needs to be frosted for the first pan. Area of a rectangle: A = (l)(w) A = (6)(9) A = 54 Surface area of a rectangular prism: SA = 2(a)(b) + 2(b)(c) + 2(a)(c) 2(6)(2) + 2(6)(9) + 2(2)(9) 2(12) + 2(54) + 2(18) 24 + 108 + 36 50 + 108 158 158 – 54 = 104 104 inches needs to be frosted. That should do it. Sorry it took so long.

OpenStudy (anonymous):

IT OK THXS

OpenStudy (anonymous):

Free internet school help... good idea

OpenStudy (anonymous):

LOL YA

OpenStudy (anonymous):

Can the numbers 24, 32, 40 be the lengths of the three sides of a right triangle? Explain why or why not. PLZZZ HELPPPP

OpenStudy (anonymous):

hmmm.... another essay? I'll try!

OpenStudy (anonymous):

OMG OK

OpenStudy (anonymous):

A^2+B^2 = C^2 24^2+32^2 = 40^2 576+1024 = 1600 1600 = 1600 Yes the side lengths can be from a right triangle. Using the Pythagorean Theorem the squares of two sides equal the square of the hypotenuse. You or your sis do FLVS? I think I remember these questions from when I did it.

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

THANK YOU VERY MUCH

OpenStudy (anonymous):

ur Welcome. I also love cookies. see ya!